# The stable trees revisited

### v1

@@ 4.2 Convergence of random finite-dimensional distributions
$\leq\sum_{l=0}^{\lfloor t_{k}m_{n}\rfloor}\frac{\widetilde{Q}_{l/m_{n}}^{n,k-1}}{m_{n}}\left(\frac{\sigma^{n}(s_{k}-\tfrac{1}{m_{n}})}{\sigma^{n}(\tfrac{l-k}{m_{n}})-\frac{m_{n}(k-1)}{n}}\wedge 1\right)\prod_{i=l+1}^{\lfloor tm_{n}\rfloor-1}\left(1-\frac{n\sigma^{n}\left(\tfrac{i-1}{m_{n}}\right)}{(n-i+k)m_{n}}\right)\frac{\sigma^{n}(t-\tfrac{1}{m_{n}})-\frac{km_{n}}{n}}{1-\frac{\lfloor tm_{n}\rfloor}{n}}$
$=\int_{0}^{t_{k}}\!\!\widetilde{Q}_{u}^{n,k-1}\!\!\left(\frac{\sigma^{n}(s_{k}-1/m_{n})}{\sigma^{n}(u-\tfrac{k}{m_{n}})-\frac{m_{n}(k-1)}{n}}\wedge 1\right)\!\prod_{i=\lfloor um_{n}\rfloor+1}^{\lfloor tm_{n}\rfloor-1}\!\left(1-\frac{n\sigma^{n}\left(\tfrac{i-1}{m_{n}}\right)}{(n-i+k)m_{n}}\right)\sigma^{n}(t-\tfrac{1}{m_{n}})\,du$
$\hskip 404.02904pt+o(1).$
@@
    $\frac{\sigma^{n}(s_{k}-1/m_{n})}{\sigma^{n}(u-k/m_{n})-\frac{m_{n}(k-1)}{n}}\wedge 1$ is uniformly close to $\frac{\sigma^{n}(u-)\wedge\sigma^{n}(s_{k})}{\sigma^{n}(u-)}$;
@@
    $\sigma^{n}(t-1/m_{n})$ is uniformly close to $\sigma^{n}(t-)$.
@@
We have already shown that, for $t_{1},\dots,t_{k}\leq T$,
@@ 4.3 Gromov–Hausdorff–Prokhorov convergence of the trees
taken at time $y_{k}^{n}$. We endow $(V(T^{n}(k)),\frac{1}{m_{n}}d^{n}|_{T^{n}(k)})$ with the empirical measure $\mu^{n}_{k}=\frac{1}{k}\sum_{i=1}^{k}\delta_{i}$ on the vertices labelled $1,2,\ldots,k$.
Analogously, we can consider the tree $\mathcal{T}$ constructed from the procedure in Section 2, and recall that $\mathcal{T}(k)=\mathcal{T}_{Y_{k}}$ is the subtree obtained by terminating the process at time $Y_{k}$. Recall also that $\mu_{k}=\frac{1}{k}\sum_{i=1}^{k}\delta_{Y_{i}}$ and that, from (8), that we have
@@
In order to complete the proof of Theorem 1.1, we need a tightness argument to deduce that $(V(T^{n}),d^{n}/m_{n},\mu^{n})\overset{d}{\to}(\mathcal{T},d,\mu)$ in the Gromov–Hausdorff–Prokhorov sense. We follow closely the strategy deployed by Haas and Miermont in Sections 4.3 and 4.4 of [haasmiermbt]. That paper proves a collection of scaling limit theorems for Markov branching trees, a family of models which includes conditioned Galton–Watson trees. Indeed, Theorem 8 (Case 2) of [haasmiermbt] is a version of Theorem 1.1 restricted to the special case of offspring distributions satisfying $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ as $k\to\infty$ for some $c>0$. Haas and Miermont’s method for showing convergence of the random finite-dimensional distributions makes use of the theory of fragmentation processes, and is completely different to ours. Their tightness proof, however, adapts well to our situation. Since we wish this paper to be self-contained, we reprove various estimates used in [haasmiermbt] via our line-breaking approach rather than appealing to the fragmentation perspective.
As argued in Section 4.4 of [haasmiermbt], by Proposition 2.4 of [EvansWinter], the convergence in distribution $(V(T^{n}),d^{n}/m_{n})\overset{d}{\to}(\mathcal{T},d)$ in the Gromov–Hausdorff sense entails that $(V(T^{n}),d^{n}/m_{n},\mu_{n})$ is a relatively compact sequence in the Gromov–Hausdorff–Prokhorov sense. Therefore, it suffices to identify any subsequential limit as $(\mathcal{T},d,\mu)$. But this then follows from the convergence of the random finite-dimensional distributions (9). Therefore, in what follows we may ignore the measures and just prove tightness for the Gromov–Hausdorff distance.
In this subsection, let us abuse notation and write $T^{n}(k)$ in place of $(V(T^{n}),d^{n}|_{V(T^{n}(k))})$ and $T^{n}$ in place of $(V(T^{n}),d^{n})$. By the principle of accompanying laws (see e.g. Theorem 3.2 of [Billingsley]), it is enough to show
@@
We adapt Lemma 33 of [haasmiermbt] to prove the following.
@@
Lemma 33 of [haasmiermbt] makes use of a hypothesis $(H^{\prime})$, which Haas and Miermont only prove holds in the restricted case $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ as $k\to\infty$ for some $c>0$. However, all that is really needed in the proof is that for fixed $\gamma$,
@@
where $A_{1},A_{2},\ldots$ are the sizes of the subtrees of the children of the root. We will give a crude bounding argument to show that this condition also holds in our case. Writing $\beta=\gamma-1>0$, we observe that
@@
The proof now follows as in Lemma 33 of [haasmiermbt]. We provide a sketch of the argument here: first, without loss of generality we can focus on large values of $q$, specifically $q>\frac{\alpha}{\alpha-1}$. Pick some $\varepsilon>0$ such that $\gamma:=q(\frac{\alpha-1}{\alpha}-\varepsilon)>1$. Now we can find a new sequence $\widetilde{m}_{n}\sim m_{n}$ such that
@@
for some fixed $N$. Now pick $M_{q}$ large enough that the following hold:
@@
We continue to proceed exactly as in [haasmiermbt]. Note that since $(m_{n})_{n\geq 1}$ is regularly varying of index $1-1/\alpha$, we have that $(m_{n}^{2\alpha/(\alpha-1)}n^{-1})_{n\geq 1}$ is regularly varying of index 1 and so we can find a constant $C$ such that
@@
for all $1\leq k\leq n$. Combining (10) and Lemma 4.8, and taking $q=2\alpha/(\alpha-1)$, we obtain that for any $n$ and $k$,
@@
For each fixed $k$, we have $F_{*}^{n}(k)/n\overset{d}{\to}F_{*}(k)$ as $n\to\infty$, where $F_{*}(k)\sim\mathrm{Beta}(1-1/\alpha,k)$ and so, in particular, $F_{*}(k)\overset{d}{\to}0$ as $k\to\infty$.
@@
$\mathbb{P}\left(F_{*}^{n}(k)\geq\lfloor xn\rfloor\ \big|\ L,N\right)=\!\!\!\sum_{j=\lfloor xn\rfloor}^{n}\!\frac{\frac{j}{n-L}N\mathbb{P}\left(F_{1}=j\ \big|\ L,N\right)\mathbb{P}\left(F_{2}+\dots+F_{N}=n-L-j\ \big|\ \!L,N\right)}{\mathbb{P}\left(F_{1}+\dots+F_{N}=n-L\ \big|\ L,N\right)}.$
@@
as $n\to\infty$, uniformly in $r$, but by Theorem 4.2.1 of [locallimits] we also have a local limit for sums of $F_{i}$: specifically, we can find a sequence $b_{n}$ such that $\frac{F_{1}+\dots+F_{n}}{b_{n}}$ converges in distribution to a $(1/\alpha)$-stable subordinator at time $1$. Letting $q$ be the density of this limit law, we have
@@
We seek a pointwise limit for $I_{k}^{n}$. Let $Y_{k}$ be the time of the $k$th cut in the continuous line-breaking construction. Then we know that
@@
  • By the local limit results for $X$ and $F$, the fraction converges to
@@
By studying the above estimates more closely, we see that there exist constants $R=R(x)\leq x^{-\alpha}+100$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and all $u\in(x,1)$ we have $I_{k}^{n}(u)\leq R\cdot\mathcal{I}_{k}(u)$. In particular, if we can show that $\int_{x}^{1}\mathcal{I}_{k}(u)du<\infty$ almost surely, then by dominated convergence it will follow that
@@
To estimate $\mathcal{I}_{k}(u)$ we can use the duality relation (14.41) in [satobook] between $p$ and $q$: we see that
@@
and as $u\to 1$ we have $\mathcal{I}_{k}(u)\to 0$ almost surely (to see this use the stretched exponential decay estimate (3) for $p$). Thus $\mathcal{I}_{k}$ is the restriction to $[x,1)$ of a continuous function on $[x,1]$ and so is bounded. We deduce that (12) holds. Now we can use the bounded convergence theorem to eliminate the conditioning, and obtain
@@
for all $x\neq 0$. If $G_{k}(0)=1$ then the convergence also holds for $x=0$; if not, then $G_{k}$ is discontinuous at 0. It follows that (13) holds at all continuity points of $G_{k}$, and so $F_{*}^{n}(k)\overset{d}{\to}F_{*}(k)$.
We now show that $F_{*}(k)\sim\mathrm{Beta}(1-1/\alpha,k)$. If this is the case then
$\mathbb{E}\left[F_{*}(k)\right]=\frac{1-1/\alpha}{k+1-1/\alpha}\to 0$
@@
Recall that $F_{*}(k)$ has density function $\phi_{k}(x)=\mathbb{E}\left[\mathcal{I}_{k}(x)\right]$. Applying the measure change at the stopping time $Y_{k}$ allows us to write this density as
$\phi_{k}(x)=x^{-\frac{1}{\alpha}}(1-x)^{-1-\frac{1}{\alpha}}\mathbb{E}\left[\sigma_{Y_{k}}\exp\left(\int_{0}^{Y_{k}}\sigma_{s}\,ds\right)p\left(-\frac{\sigma_{Y_{k}}}{(1-x)^{\frac{1}{\alpha}}}\right)\right].$
@@
$\phi_{k}(x)=\frac{x^{-\frac{1}{\alpha}}(1-x)^{-1-\frac{1}{\alpha}}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{t}^{2}\left(\int_{0}^{t}\sigma_{s}\,ds\right)^{k-1}p\left(-\frac{\sigma_{t}}{(1-x)^{\frac{1}{\alpha}}}\right)\right]\,dt.$
@@
$\frac{x^{-\frac{1}{\alpha}}(1-x)^{-1+\frac{1}{\alpha}}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left((1-x)^{\frac{1}{\alpha}}\int_{0}^{t}\sigma_{s(1-x)^{-1+\frac{1}{\alpha}}}\,ds\right)^{k-1}\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{x^{-\frac{1}{\alpha}}(1-x)^{-1+\frac{1}{\alpha}}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left((1-x)\int_{0}^{t(1-x)^{-1+\frac{1}{\alpha}}}\sigma_{r}\,dr\right)^{k-1}\!\!\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{x^{-\frac{1}{\alpha}}(1-x)^{k-2+\frac{1}{\alpha}}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left(\int_{0}^{t(1-x)^{-1+\frac{1}{\alpha}}}\sigma_{r}\,dr\right)^{k-1}\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{x^{-\frac{1}{\alpha}}(1-x)^{k-1}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{u}^{2}\left(\int_{0}^{u}\sigma_{r}\,dr\right)^{k-1}\!\!\!p\left(-\sigma_{u}\right)\right]\,du,$
after various changes of variable. The integral term in the last expression does not depend on $x$ and thus acts only as a normalising constant for the density. In particular, we can identify this density as that of the $\text{Beta}(1-\frac{1}{\alpha},k)$ distribution, and it also follows that
$\int_{0}^{\infty}\mathbb{E}\left[\sigma_{u}^{2}\left(\int_{0}^{u}\sigma_{r}dr\right)^{k-1}p(-\sigma_{u})\right]du=\frac{\Gamma(k+1-1/\alpha)}{\Gamma(1-1/\alpha)}.\qed$
@@
$F_{*}(k)$ may clearly be interpreted as a size-biased pick from among the component sizes of the forest $\mathcal{T}\setminus\mathcal{T}(k)$. To the best of our knowledge, its distribution had not previously been explicitly identified in the literature for a general $k$, although it is easy to deduce it from pre-existing results, as we now show. In the case $k=1$, the ranked sequence of component sizes of $\mathcal{T}\setminus\mathcal{T}(1)$ forms the fine spinal mass partition defined in [spinalpartition]. By Corollary 10 of the same paper, the sequence of sizes follows a Poisson–Dirichlet distribution with parameters $(\frac{1}{\alpha},1-\frac{1}{\alpha})$. (See Chapter 3 of Pitman [PitmanCSP] for more information about the Poisson–Dirichlet distributions.)
Now let $P^{k}=(P_{1}^{k},P_{2}^{k},\dots)$ be the ranked sequence of component sizes of $\mathcal{T}\setminus\mathcal{T}(k)$. Conditionally given $P^{k-1}$, we can sample $P^{k}$ as follows: we take a size-biased pick from the components of $\mathcal{T}\setminus\mathcal{T}(k-1)$, then reveal the fine spinal mass partition of this component. This has the effect of replacing some $P_{I}^{k-1}$ with the collection $\{P_{I}^{k-1}\tilde{P}_{i}:i\in\mathbb{N}\}$, where $I$ is a size-biased index, $\tilde{P}$ is a copy of $P^{1}$ independent of everything else, and finally we reorder the sequence to preserve the decreasing order. By Theorem 3.1 in [repeatedspinalpartitionlaw], we see that if $P^{k-1}\sim\mathrm{PD}(\frac{1}{\alpha},\theta)$ for some $\theta>0$, then $P^{k}\sim\mathrm{PD}(\frac{1}{\alpha},\theta+1)$. It follows by induction that $P^{k}\sim\mathrm{PD}(\frac{1}{\alpha},k-\frac{1}{\alpha})$ for every $k\geq 1$. Now $F_{*}(k)$ is just a size-biased pick from $P^{k}$, and thus it has law $\text{Beta}(1-\frac{1}{\alpha},k)$.

### v2

@@ 4.2 Convergence of random finite-dimensional distributions
$\leq\sum_{l=0}^{\lfloor t_{k}m_{n}\rfloor}\frac{\widetilde{Q}_{l/m_{n}}^{n,k-1}}{m_{n}}\!\left(\frac{\sigma^{n}(s_{k})}{\sigma^{n}(\tfrac{l-k}{m_{n}})-\frac{m_{n}(k-1)}{n}}\wedge 1\right)\!\!\prod_{i=l+1}^{\lfloor tm_{n}\rfloor-1}\!\!\left(\!1-\frac{n\sigma^{n}\left(\tfrac{i-1}{m_{n}}\right)}{(n-i+k)m_{n}}\!\right)\!\!\frac{\sigma^{n}(t-\tfrac{k+1}{m_{n}})-\frac{km_{n}}{n}}{1-\frac{\lfloor tm_{n}\rfloor}{n}}$
$=\int_{0}^{t_{k}}\!\!\widetilde{Q}_{u}^{n,k-1}\!\!\left(\frac{\sigma^{n}(s_{k})}{\sigma^{n}(u-\tfrac{k}{m_{n}})-\frac{m_{n}(k-1)}{n}}\wedge 1\right)\!\prod_{i=\lfloor um_{n}\rfloor+1}^{\lfloor tm_{n}\rfloor-1}\!\left(1-\frac{n\sigma^{n}\left(\tfrac{i-1}{m_{n}}\right)}{(n-i+k)m_{n}}\right)\sigma^{n}(t-\tfrac{1}{m_{n}})\,du$
$\hskip 355.65944pt+o(1).$
@@
    $\frac{\sigma^{n}(s_{k})}{\sigma^{n}(u-k/m_{n})-\frac{m_{n}(k-1)}{n}}\wedge 1$ is uniformly close to $\frac{\sigma^{n}(u-)\wedge\sigma^{n}(s_{k})}{\sigma^{n}(u-)}$;
@@
    $\sigma^{n}(t-(k+1)/m_{n})$ is uniformly close to $\sigma^{n}(t-)$.
@@
We have already shown that, for $t_{1},\dots,t_{k}\leq T$ and $s_{1},\dots,s_{k-1}$ with $s_{i}\leq t_{i}$,
@@ 4.3 Gromov–Hausdorff–Prokhorov convergence of the trees
taken at time $y_{k}^{n}$. We endow $(V(T^{n}(k)),\frac{1}{m_{n}}d^{n}|_{V(T^{n}(k))})$ with the empirical measure $\mu^{n}_{k}=\frac{1}{k}\sum_{i=1}^{k}\delta_{i}$ on the vertices labelled $1,2,\ldots,k$.
Analogously, we can consider the tree $\mathcal{T}$ constructed from the procedure in Section 2, and recall that $\mathcal{T}(k)=\mathcal{T}_{Y_{k}}$ is the subtree obtained by terminating the process at time $Y_{k}$. Recall also that $\mu_{k}=\frac{1}{k}\sum_{i=1}^{k}\delta_{Y_{i}}$ and that, from (8), we have
@@
In order to complete the proof of Theorem 1.1, we need a tightness argument to deduce that $(V(T^{n}),d^{n}/m_{n},\mu^{n})\overset{d}{\to}(\mathcal{T},d,\mu)$ in the Gromov–Hausdorff–Prokhorov sense. We follow closely the strategy deployed by Haas and Miermont in Sections 4.3 and 4.4 of [28]. That paper proves a collection of scaling limit theorems for Markov branching trees, a family of models which includes conditioned Galton–Watson trees. Indeed, Theorem 8 (Case 2) of [28] is a version of Theorem 1.1 restricted to the special case of offspring distributions satisfying $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ as $k\to\infty$ for some $c>0$. Haas and Miermont’s method for showing convergence of the random finite-dimensional distributions makes use of the theory of fragmentation processes, and is completely different to ours. Their tightness proof, however, adapts well to our situation. Since we wish this paper to be self-contained, we reprove various estimates used in [28] via our line-breaking approach rather than appealing to the fragmentation perspective.
As argued in Section 4.4 of [28], by Proposition 2.4 of [26], the convergence in distribution $(V(T^{n}),d^{n}/m_{n})\overset{d}{\to}(\mathcal{T},d)$ in the Gromov–Hausdorff sense entails that $(V(T^{n}),d^{n}/m_{n},\mu_{n})$ is a relatively compact sequence in the Gromov–Hausdorff–Prokhorov sense. Therefore, it suffices to identify any subsequential limit as $(\mathcal{T},d,\mu)$. But this then follows from the convergence of the random finite-dimensional distributions (9). Therefore, in what follows we may ignore the measures and just prove tightness for the Gromov–Hausdorff distance.
In this subsection, let us abuse notation and write $T^{n}$ in place of $(V(T^{n}),d^{n})$ and $T^{n}(k)$ in place of $(V(T^{n}(k)),d^{n}|_{V(T^{n}(k))})$. By the principle of accompanying laws (see e.g. Theorem 3.2 of [14]), it is enough to show
@@
We adapt Lemma 33 of [28] to prove the following.
@@
Lemma 33 of [28] makes use of a hypothesis $(H^{\prime})$, which Haas and Miermont only prove holds in the restricted case $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ as $k\to\infty$ for some $c>0$. However, all that is really needed in the proof is that for fixed $\gamma>1$,
@@
where $A_{1},A_{2},\ldots$ are the sizes of the subtrees from the children of the root. We will give a crude bounding argument to show that this condition also holds in our case. Writing $\beta=\gamma-1>0$, we observe that
@@
The proof now follows as in Lemma 33 of [28]. We provide a sketch of the argument here: first, without loss of generality we can focus on large values of $q$, specifically $q>\frac{\alpha}{\alpha-1}$. Pick some $\varepsilon>0$ such that $\gamma:=q(\frac{\alpha-1}{\alpha}-\varepsilon)>1$. Now we can find a new sequence $\widetilde{m}_{n}\sim m_{n}$ such that
@@
for some fixed $\tilde{M}$ and $N$. Now pick $M_{q}$ large enough that the following hold:
@@
We continue to proceed exactly as in [28]. Note that since $(m_{n})_{n\geq 1}$ is regularly varying of index $1-1/\alpha$, we have that $(m_{n}^{2\alpha/(\alpha-1)}n^{-1})_{n\geq 1}$ is regularly varying of index 1 and so we can find a constant $C$ such that
@@
for all $1\leq k\leq n$. Combining (10) and 4.8, and taking $q=2\alpha/(\alpha-1)$, we obtain that for any $n$ and $k$,
@@
For each fixed $k$, we have $F_{*}^{n}(k)/n\overset{d}{\to}F_{*}(k)$ as $n\to\infty$, where $F_{*}(k)$ is a random variable stochastically dominated by the $\mathrm{Beta}(1-1/\alpha,k)$ distribution and so, in particular, $F_{*}(k)\overset{d}{\to}0$ as $k\to\infty$.
@@
$\mathbb{P}\left(F_{*}^{n}(k)\geq\lfloor xn\rfloor\ \big|\ L,N\right)$
$\qquad=\sum_{j=\lfloor xn\rfloor}^{n}\!\frac{\frac{j}{n-L}N\mathbb{P}\left(F_{1}=j\ \big|\ L,N\right)\mathbb{P}\left(F_{2}+\dots+F_{N}=n-L-j\ \big|\ \!L,N\right)}{\mathbb{P}\left(F_{1}+\dots+F_{N}=n-L\ \big|\ L,N\right)}.$
@@
as $n\to\infty$, uniformly in $r$, but by Theorem 4.2.1 of [31] we also have a local limit for sums of $F_{i}$: specifically, we can find a sequence $b_{n}$ such that $\frac{F_{1}+\dots+F_{n}}{b_{n}}$ converges in distribution to a $(1/\alpha)$-stable subordinator at time $1$. Letting $q$ be the density of this limit law, we have
@@
We seek an almost sure pointwise limit for $I_{k}^{n}$. Let $Y_{k}$ be the time of the $k$th cut in the continuous line-breaking construction. Then we know that
@@
  • By the local limit results for $S$ and $F$, the fraction converges to
@@
By studying the above estimates more closely, we see that there exist constants $R=R(x)\leq x^{-\alpha}+100$ and (random) $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and all $u\in(x,1)$ we have $I_{k}^{n}(u)\leq R\cdot\mathcal{I}_{k}(u)$. In particular, if we can show that $\int_{x}^{1}\mathcal{I}_{k}(u)du<\infty$ almost surely, then by dominated convergence it will follow that
@@
To estimate $\mathcal{I}_{k}(u)$ we can use the duality relation (14.41) in [39] between $p$ and $q$: we see that
@@
and as $u\to 1$ we have $\mathcal{I}_{k}(u)\to 0$ almost surely (to see this use the stretched exponential decay estimate (3) for $p$). Thus $\mathcal{I}_{k}$ is almost surely the restriction to $[x,1)$ of a continuous function on $[x,1]$ and so is bounded. We deduce that (12) holds. Now we can use the bounded convergence theorem to eliminate the conditioning, and obtain
@@
for all $x\neq 0$. If $G_{k}(0)=1$ then the convergence also holds for $x=0$; if not, then $G_{k}$ is discontinuous at 0. It follows that (13) holds at all continuity points of $G_{k}$, and so $F_{*}^{n}(k)/n\overset{d}{\to}F_{*}(k)$.
It remains to identify the law of $F_{*}(k)$. We will show that it is a mixture of a point mass at 0 (with weight $1-G_{k}(0)$) and a $\text{Beta}(1-\frac{1}{\alpha},k)$ distribution (with weight $G_{k}(0)$). If this is the case then
$\mathbb{E}\left[F_{*}(k)\right]=G_{k}(0)\cdot\frac{1-1/\alpha}{k+1-1/\alpha}\leq\frac{1-1/\alpha}{k+1-\frac{1}{\alpha}}\to 0$
@@
We study the conditional law of $F_{*}(k)$ given that $F_{*}(k)>0$. This has density function $\phi_{k}(x)=c\mathbb{E}\left[\mathcal{I}_{k}(x)\right]$, where $c=1/G_{k}(t)$. Applying the measure change at the stopping time $Y_{k}$ allows us to write this density as
$\phi_{k}(x)=cx^{-\frac{1}{\alpha}}(1-x)^{-1-\frac{1}{\alpha}}\mathbb{E}\left[\sigma_{Y_{k}}\exp\left(\int_{0}^{Y_{k}}\sigma_{s}\,ds\right)p\left(-\frac{\sigma_{Y_{k}}}{(1-x)^{\frac{1}{\alpha}}}\right)\right].$
@@
$\phi_{k}(x)=\frac{cx^{-\frac{1}{\alpha}}(1-x)^{-1-\frac{1}{\alpha}}}{(k-1)!}\int_{0}^{\infty}\mathbb{E}\left[\sigma_{t}^{2}\left(\int_{0}^{t}\sigma_{s}\,ds\right)^{k-1}p\left(-\frac{\sigma_{t}}{(1-x)^{\frac{1}{\alpha}}}\right)\right]\,dt.$
@@
$\frac{cx^{-\frac{1}{\alpha}}(1-x)^{-1+\frac{1}{\alpha}}}{(k-1)!}\!\!\int_{0}^{\infty}\!\!\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left((1-x)^{\frac{1}{\alpha}}\int_{0}^{t}\!\sigma_{s(1-x)^{-1+\frac{1}{\alpha}}}\,ds\right)^{k-1}\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{cx^{-\frac{1}{\alpha}}(1-x)^{-1+\frac{1}{\alpha}}}{(k-1)!}\!\!\int_{0}^{\infty}\!\!\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left((1-x)\int_{0}^{t(1-x)^{-1+\frac{1}{\alpha}}}\!\!\!\!\!\sigma_{r}\,dr\right)^{k-1}\!\!\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{cx^{-\frac{1}{\alpha}}(1-x)^{k-2+\frac{1}{\alpha}}}{(k-1)!}\!\!\int_{0}^{\infty}\!\!\mathbb{E}\left[\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}^{2}\left(\int_{0}^{t(1-x)^{-1+\frac{1}{\alpha}}}\sigma_{r}\,dr\right)^{k-1}\!\!\!p\left(-\sigma_{t(1-x)^{-1+\frac{1}{\alpha}}}\right)\right]\,dt$
$=\frac{x^{-\frac{1}{\alpha}}(1-x)^{k-1}}{(k-1)!}\cdot c\int_{0}^{\infty}\mathbb{E}\left[\sigma_{u}^{2}\left(\int_{0}^{u}\sigma_{r}\,dr\right)^{k-1}\!\!\!p\left(-\sigma_{u}\right)\right]\,du,$
after various changes of variable. The integral term in the last expression does not depend on $x$ and thus acts only as a normalising constant for the density. In particular, we can identify this density as that of the $\text{Beta}(1-\frac{1}{\alpha},k)$ distribution.∎
@@
We expect that, in fact, $F_{*}(k)$ has no mass at zero and has precisely a $\text{Beta}(1-\frac{1}{\alpha},k)$ distribution, and should be interpreted as a size-biased pick from among the component sizes of the forest $\mathcal{T}\setminus\mathcal{T}(k)$. Indeed such a size-biased pick has distribution $\text{Beta}(1-\frac{1}{\alpha},k)$. To the best of our knowledge, this has not previously been explicitly identified in the literature for a general $k$, although it is easy to deduce it from pre-existing results, as we now show. In the case $k=1$, the ranked sequence of component sizes of $\mathcal{T}\setminus\mathcal{T}(1)$ forms the fine spinal mass partition defined in [29]. By Corollary 10 of the same paper, the sequence of sizes follows a Poisson–Dirichlet distribution with parameters $(\frac{1}{\alpha},1-\frac{1}{\alpha})$. (See Chapter 3 of Pitman [37] for more information about the Poisson–Dirichlet distributions.)
Now let $P^{k}=(P_{1}^{k},P_{2}^{k},\dots)$ be the ranked sequence of component sizes of $\mathcal{T}\setminus\mathcal{T}(k)$. Conditionally given $P^{k-1}$, we can sample $P^{k}$ as follows: we take a size-biased pick from the components of $\mathcal{T}\setminus\mathcal{T}(k-1)$, then reveal the fine spinal mass partition of this component. This has the effect of replacing some $P_{I}^{k-1}$ with the collection $\{P_{I}^{k-1}\tilde{P}_{i}:i\in\mathbb{N}\}$, where $I$ is a size-biased index, $\tilde{P}$ is a copy of $P^{1}$ independent of everything else, and finally we reorder the sequence to preserve the decreasing order. By Theorem 3.1 in [21], we see that if $P^{k-1}\sim\mathrm{PD}(\frac{1}{\alpha},\theta)$ for some $\theta>0$, then $P^{k}\sim\mathrm{PD}(\frac{1}{\alpha},\theta+1)$. It follows by induction that $P^{k}\sim\mathrm{PD}(\frac{1}{\alpha},k-\frac{1}{\alpha})$ for every $k\geq 1$. Now $F_{*}(k)$ is just a size-biased pick from $P^{k}$, and thus it has law $\text{Beta}(1-\frac{1}{\alpha},k)$.